OpenStudy (anonymous):

A 10.0 Ω resistor is in parallel with a 20.0 Ω resistor and supplied by a 9.00 V battery. What is the total current passing through the circuit? A. 0.38 A B. 1.35 A C. 1.13 A D. 6.00 A

OpenStudy (michele_laino):

the total or equivalent resistance is: \[\Large {R_{EQ}} = \frac{1}{{\frac{1}{{10}} + \frac{1}{{20}}}} = ...\Omega \]

OpenStudy (anonymous):

6.6666666?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so the total current is: I=90/6.67=...amperes

OpenStudy (anonymous):

13.49?

OpenStudy (michele_laino):

ok! More precisely it is 13.5 amperes

OpenStudy (anonymous):

ohh my choices gives me 1.35 A.. would that be the solutoin?

OpenStudy (michele_laino):

sorry, it is: 9/6.67 which is equal to 1.35 amperes

OpenStudy (anonymous):

oh okay! yay! thanks!!:D

OpenStudy (michele_laino):

:)